OpenStudy (madgirlwithabluebox):

Will medal. Solve the system by substitution. y=x-1 y=-2x+5

OpenStudy (anonymous):

put the value of y in y=-2x+5

OpenStudy (madgirlwithabluebox):

so x-1=-2x+5?

OpenStudy (anonymous):

the answer is y=-7

OpenStudy (madgirlwithabluebox):

thats nice and thanks but i really wanted to know how to do it on my own

OpenStudy (anonymous):

sry :(

OpenStudy (madgirlwithabluebox):

could you show me how you got y=-7?

OpenStudy (madgirlwithabluebox):

@surjithayer ?

OpenStudy (anonymous):

x-1=-2x+5 x+2x=5+1 3x=6 x=6/3=? then y=x-1 or y=?-1=?